OpenStudy (anonymous):

how do i find the values of x when i have cos^2x = 2sinx?

OpenStudy (jdoe0001):

hmmm do you know the trig identity of \(\bf sin^2(\theta)+cos^2(\theta)=?\)

OpenStudy (anonymous):

it = 1

OpenStudy (jdoe0001):

so... if we solve that equation for say \(\bf cos^2(\theta)\) what would that give us?

OpenStudy (anonymous):

1-sin^2x

OpenStudy (jdoe0001):

ohhh .. ok . so \(\bf sin^2(\theta)+cos^2(\theta)=1\implies {\color{blue}{ cos^2(\theta)}}=1-sin^2(\theta)\qquad thus\\ \quad \\ {\color{blue}{ cos^2(x)}} = 2sin(x)\implies {\color{blue}{ 1-sin^2(x)}}=2sin(x)\\ \quad \\ 0=sin^2(x)+2sin(x)-1\) and as you can see, now you're just left with a quadratic, so just solve for \(\bf sin(x)\)

OpenStudy (anonymous):

how would i solve the problem?

OpenStudy (anonymous):

how do i get rid of the sinx on both sides of the equation?

OpenStudy (anonymous):

its a quadratic equation now do you see it?

OpenStudy (jdoe0001):

well, think about it just like you'd be solving say \(\bf x^2+2x-1\)

OpenStudy (anonymous):

i tried doing the 2 pairs of parantheses but that does not work

OpenStudy (jdoe0001):

hmm then use the quadratic formula

OpenStudy (jdoe0001):

\(\bf 0=sin^2(x)+2sin(x)-1\\ \quad \\ sin(x)= \cfrac{ - 2 \pm \sqrt { 2^2 -4(1)(-1)}}{2(1)}\)

OpenStudy (jdoe0001):

\(\bf 0=sin^2(x)+2sin(x)-1\\ \quad \\ sin(x)= \cfrac{ - 2 \pm \sqrt { 2^2 -4(1)(-1)}}{2(1)}\\ \quad \\ sin(x)= \cfrac{ - 2 \pm \sqrt { 4+4}}{2(1)}\implies sin(x)= \cfrac{ - 2 \pm \sqrt { 8}}{2(1)}\\ \quad \\ sin(x)= \cfrac{ - \cancel{2} \pm \cancel{2}\sqrt { 2}}{\cancel{2}}\)

OpenStudy (anonymous):

yes i got the same answer

OpenStudy (jdoe0001):

\(\bf sin(x)= \cfrac{ - 2 \pm 2\sqrt { 2}}{2(1)}\quad \textit{taking }sin^{-1}\\ \quad \\ sin^{-1}[sin(x)]= sin^{-1}\left[-1\pm\sqrt{2}\right]\implies x=sin^{-1}\left[-1\pm\sqrt{2}\right]\)

OpenStudy (jdoe0001):

\(\bf sin(x)= -1\pm\sqrt{2}\\ \quad \\ sin^{-1}[sin(x)]= sin^{-1}\left[-1\pm\sqrt{2}\right]\implies x=sin^{-1}\left[-1\pm\sqrt{2}\right]\) for that matter... anyhow, that'd give the values for "x"

OpenStudy (anonymous):

i got Error on my calculator for one of the 2 x values

OpenStudy (anonymous):

the other x value was about 24. 47

OpenStudy (jdoe0001):

right, the negative root, because the -1 will add to it, so that'd be an extraneous value, so you'd discard that one it gives -2.41 which is a value outside the sine function range, -1 < sine < 1 so the valid one is the positive root one

OpenStudy (jdoe0001):

well \(\bf -1 \le sin(x) \le +1\)

OpenStudy (jdoe0001):

he other x value was about 24. 47 \(\large \checkmark\)